OpenStudy (anonymous):

Solve x(dy/dx)=y+2xe^(-y/x) I'm stuck (FO nonlinear DE)

OpenStudy (anonymous):

are you saying derive in respect to x?

OpenStudy (anonymous):

or is this a multivariable problem?

OpenStudy (anonymous):

\[x dy/dx = y+2xe^(-y/x)\]

OpenStudy (anonymous):

its suppose to be a First Order Nonlinear problem... my teacher wants the equation that solve it

OpenStudy (anonymous):

It can be solved by separating variables.

OpenStudy (anonymous):

Its a separable equation?

OpenStudy (anonymous):

I think so.

OpenStudy (anonymous):

I tried separating but it didn't look pretty lol

OpenStudy (anonymous):

e^(-y/x)=e^(-y).e^(1/x)

OpenStudy (anonymous):

If that helps.

OpenStudy (anonymous):

i thought e^(-y) * e^(1/x) = e^(-y+(1/x)) ?

OpenStudy (anonymous):

:( Something is terribly wrong with me.

OpenStudy (anonymous):

You're right :) I'll try another method.

OpenStudy (anonymous):

thanks, i tried that already... something was wrong with me too lol

OpenStudy (anonymous):

I'm so lost on this one...